# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### v1

@@ 12.4 Phase space path integral for quark loop amplitudes
$\int_{0}^{\infty}de\exp\left(-ed\theta(\mathrm{i}\gamma_{\mu}P_{\mu}+m_{q})\right)\propto\frac{1}{(\mathrm{i}\gamma_{\mu}P_{\mu}(\theta)+m_{q})}$ (12.12)
@@ 12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
This reduction from $4\times 4$ to $2\times 2$ matrices significantly reduces the computational complexity for numerical evaluation. The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates. Therefore, the Complex Langevin method [42] required to compute this path integral.
@@ 13 The Linear Regge trajectories in the WKB limit
13.1 The Helicoid Saddle Point
In the semiclassical limit (large momentum or large tension $\sigma$), the path integral is dominated by the saddle point of the effective action. We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
@@ 13.1 The Helicoid Saddle Point
In the spinor formalism, with rescaling $w(z)=z^{-1/2}$ the Helicoid is generated by:
$\lambda(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}1\\ -z\end{pmatrix}$ (13.2)
Substituting this into the null vector definition $f^{\prime}_{\mu}\sigma_{\mu}=\lambda\otimes\mu$, we find:
$f^{\prime}_{1}+\mathrm{i}f^{\prime}_{2}\sim\frac{1}{z},\qquad f^{\prime}_{3}=f^{\prime}_{4}=0.$ (13.3)
Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid. The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
13.2 The Minkowski Helicoid
@@
The induced Minkowski metric is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
@@
Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space. Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity). By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space. Thus, the Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
@@ 13.3 Semiclassical Motivation: Twisted Boundary Conditions
Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$. In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor333the perimeter renormalization terms in that factor are included in the physical quark mass. This is relevant to our WKB approximation in the Wilson loop at large loops:
@@ 13.4 Effective Lagrangean
We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks. The confining factor introduces an effective Minkowski area term. We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m\int ds$. The total Lagrangian is:
@@
The Helicoid contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is:
$L_{\text{helicoid}}$ $=-\sigma\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
$=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (13.9)
Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional. Adding the particle term $2m\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
$L(v,\omega)=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (13.10)
It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space. Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$. We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which , as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\kappa$.
@@
$J=\frac{\sigma}{\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2mv^{2}}{\omega\sqrt{1-v^{2}}}$ (13.12)
@@
$E=\frac{2\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (13.13)
@@ 13.5 Balance of quark forces and boundary condition
Recall that the total Lagrangian $L$ is composed of the bulk integral for the helicoid and the boundary terms for the quarks:
$L=-2\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (13.14)
The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$. Differentiating the integral using the Leibniz rule yields the tension force, while differentiating the boundary term yields the centrifugal force:
$\frac{\partial L}{\partial R}=-2\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (13.15)
Substituting $v=\omega R$ and computing the derivative of the mass term:
$-2\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-\omega^{2}R^{2}}}\right)=0$ (13.16)
@@
$\sigma\sqrt{1-v^{2}}=\frac{m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+m_{q}}$ (13.17)
This equation balances the confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
@@
These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$. To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$). In this limit, the mass terms vanish, and the system is governed purely by the confining geometry:
$E\approx\frac{2\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{\omega^{2}}\frac{\pi}{2}$ (13.19)
@@
$J=\frac{\sigma(\pi/2)}{(\pi\sigma/E)^{2}}\cdot\frac{\pi}{2}\implies J=\frac{E^{2}}{2\pi\sigma}$ (13.20)
@@
We thus recover the linear Regge trajectory:
$E^{2}=2\pi\sigma J$ (13.21)

### v2

@@ 12.4 Phase space path integral for quark loop amplitudes
[…]
@@ 12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
This reduction from $4\times 4$ to $2\times 2$ matrices significantly reduces the computational complexity for numerical evaluation. The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates. It is precisely this highly oscillatory nature of the path integral that mandates its exact analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
13 Twistor Holography and the Rigid Twistor String
13.1 Twistor Holography (Gauge Holography)
13.2 Analytic Twistor String vs. Topological Twistor String
Because the path integral over the twistor target space is highly oscillatory, one might assume that extracting the physical spectrum requires massive stochastic lattice simulations. However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted exactly and analytically. Planar QCD ceases to be a theory of stochastic quantum fluctuations; it is a theory of pure classical complex geometry.
14 Analytic continuation of Twistor String Path integral
The Euclidean path integral of Twistor String theory, as formulated so far, is similar to a partition function of a statistical system with positive definite energy , given by the Liouville bulk integral. There are, however, some pre-exponential factors that violate this statistical interpretation. First, there is a "vertex operator" factor
[…]
Second, there is a trace factor (12.30). Both of these factors are not positive, and they oscillate as functionals of our dynamical variables: i.e., the complex curve $\lambda(z),\mu(z)$ in the Twistor space $\mathcal{C}P^{3}$. These oscillations become even more obvious when the path integral is analytically continued in Minkowski space, where the effective action is imaginary, and the exponential decay of the statistical weight becomes a quantum oscillation of the Feynman-Dirac weight $\exp\left(\mathrm{i}S/\bar{h}\right)$.
In ordinary quantum mechanics, this phenomenon is well known: the path integral, when formulated in terms of physical, real trajectories in phase space, is exponentially small as a result of the interference of the classical trajectories in the Feynman path integral. However, when reformulated as a complex path integral by analytic continuation, it is dominated by complex saddle points, where there is no interference along the steepest descent path. The spectrum of certain quantum mechanical systems can be exactly computed by finding the dominant complex trajectory winding around the complex branch point of the analytically continued effective Action. This is a mechanism for the conversion of the Bohr-Sommerfeld quantization condition from the WKB approximation to an exact equation for the spectrum [6, 19, 9].
We are looking for a similar phenomenon in our Twistor String theory, and the first step towards such a solution is the analytic continuation of the path integral into the complex plane in parameter space.
14.1 Complexification of the parameter space
To ensure the strict analyticity of the path integral, we must remove complex conjugation from the effective action and its parameter space. We achieve this by complexifying the target manifold and analytically continuing the action integral itself. This ensures that the drift forces are partial derivatives of a globally holomorphic function. Then the path integral can be regarded as a complex integral in multi-dimensional space of holomorphic functions in Twistor space.
We replace the original real and Hermitian variables with a set of independent holomorphic variables. The angular coordinates $\alpha_{k}$ are mapped to the complex plane via $z_{1}=e^{i\alpha_{1}},\;z_{2}=e^{i\alpha_{2}}$. The spinor fields and their conjugates are treated as independent complex vectors. The full set of dynamical variables for the discretized loop is:
$\mathcal{V}=\{\{z_{1},z_{2}\},\{\rho_{k},\xi_{k},\eta_{k},\tilde{\xi}_{k},\tilde{\eta}_{k},\gamma_{1,k},\gamma_{2,k}\}_{k=1}^{N}\}.$ (14.1)
$S_{\text{constr}}=i\sum_{k}\left[\gamma_{1,k}(\tilde{\xi}_{k}\xi_{k}-1)+\gamma_{2,k}(\tilde{\eta}_{k}\eta_{k}-1)\right].$ (14.2)
14.2 Spectral Interpolation of Fields
14.2.1 Spinor Fields
$\Lambda(z)=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}K\left(ze^{-\mathrm{i}\theta_{k}}\right);$ (14.3)
$K(z)=\frac{1-z^{N}}{1-z};$ (14.4)
$\lambda(z)=\sum_{n=0}^{N-1}L_{n}z^{n},\qquad\tilde{\lambda}(z)=\sum_{n=0}^{N-1}\tilde{L}_{n}z^{n}.$ (14.5)
$\{L_{n},M_{n}\}=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}e^{-\mathrm{i}\theta_{k}}$ (14.6)
Remark 14.1.
Reduction to Pole Moduli and the Algebraic Jacobian
While the path integral is formulated using the discrete boundary values at the roots of unity $\theta_{k}=2\pi k/N$, the physical saddle points and their monodromies are governed by the meromorphic structure of the twistors—specifically, the positions $z_{i}$ and residues $R_{i}$ of their poles inside the unit disk.The transformation of the integration measure from the boundary values to these pole moduli is exact and purely algebraic. For a twistor field parameterized by a finite number of poles, its evaluation on the boundary circle $z=e^{i\theta}$ yields a geometric series:
$\lambda(e^{i\theta})=\sum_{i}\frac{R_{i}}{e^{\mathrm{i}\theta}-z_{i}}=-\sum_{i}\frac{R_{i}}{z_{i}}\sum_{n=0}^{\infty}z_{i}^{-n}e^{\mathrm{i}n\theta}$ (14.7)
Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are exactly determined as algebraic superpositions of the pole parameters:
$L_{n}=-\sum_{i}\frac{R_{i}}{z_{i}^{n+1}}$ (14.8)
Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves exactly two finite, calculable matrix Jacobians:1. The constant Jacobian of the Discrete Fourier Transform matrix (the Fourier matrix $e^{2\pi\mathrm{i}kn/N}$), which maps the boundary values to the Fourier coefficients $L_{n}$.2. The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute strictly to the pre-exponential fluctuation determinant of the path integral. For the extraction of the exact geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action. As we will demonstrate, the S-matrix poles emerge from the constructive interference of the geometric series over the topological winding number $k$. In this resonant limit, the multi-valued action scales as $k\Delta S_{eff}$. Consequently, the transverse quantum fluctuations around the saddle point are dynamically damped by a factor of $1/k$. This perfectly crushes the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles. However, the explicit existence of this algebraic mapping mathematically guarantees that our twistor measure is strictly finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
14.2.2 Liouville Field
$\rho(z,\bar{z})=\frac{\log\tilde{\lambda}(\bar{z})\lambda(z)+\log\tilde{\mu}(\bar{z})\mu(z)}{2}$ (14.9)
where $\lambda(z),\mu(z)$} is extended by interpolation (14.3). The derivatives of the above kernel $K^{\prime}(z)$ involved in the Liouville Lagrangian with this $\rho$ were computed in the previous sections.
14.3 The Boundary Terms in the action
[…]
$G_{p}(z_{1},z_{2})\equiv\int_{z_{1}}^{z_{2}}z^{p}\frac{dz}{iz}=\frac{z_{2}^{p}-z_{1}^{p}}{ip}.$ (14.10)
$S_{q}=-\mathrm{i}q_{\alpha}$
$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}\sigma_{\alpha}M_{m})-(\tilde{L}_{n}\sigma^{\dagger}_{\alpha}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (14.11)
$S_{\gamma}=\mathrm{i}\sum_{k=0}^{N-1}\gamma_{1,k}\left(\tilde{\xi}_{k}\xi_{k}-1\right)+\gamma_{2,k}\left(\tilde{\eta}_{k}\eta_{k}-1\right)$ (14.12)
[…]
$\mathcal{P}(\theta)=2\mathrm{i}\left[z\eta\tilde{\xi}^{T}-\bar{z}\xi\eta^{T}\right]_{z=e^{\mathrm{i}\theta}}$ (14.15)
Finally, our reweighting adds one more complex term to the action,
$S_{O}=\log\mathcal{O};$ (14.16)
$\mathcal{O}=\frac{\partial S_{q}}{\partial q_{4}}=$
$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (14.17)
14.4 Stochastic Quantization and Complex Langevin Dynamics
The exact mathematical evaluation of this highly oscillatory complexified path integral is rigorously achieved via Stochastic Quantization. The quantum statistical evolution of the twistor string is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{total}=S_{q}+S_{\gamma}+S_{T}+S_{\rho}+S_{\mathcal{O}}$. We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an exact continuous mathematical reformulation of the quantum path integral.
  1. 1.
    Coordinate Evolution ($z_{k}$): The drift force on the boundary points $z_{k}$ arises from the boundary terms of the $G$ functions in $S_{q}$ and the integration limits in $S_{bulk}$:
    $\frac{dz_{k}}{d\tau}=-z_{k}^{2}\frac{\partial S_{total}}{\partial z_{k}}+\mathrm{i}z_{k}\nu_{z_{k}}$ (14.18)
    (The factor $z_{k}^{2}$ arises from the metric of the map $\alpha=-\mathrm{i}\log z$, and $\nu$ represents complex Gaussian white noise).
  2. 2.
    Spinor Evolution ($\xi_{k},\tilde{\xi}_{k}$, etc.): The drift forces are computed via the chain rule through the spectral coefficients $L_{n}$. For example:
    $\frac{d\xi_{k}}{d\tau}=-\frac{\partial S_{total}}{\partial\xi_{k}}-\mathrm{i}\gamma_{1,k}\tilde{\xi}_{k}+\nu_{\xi_{k}}$ (14.19)
    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the exact quantum fluctuations of the twistor fields.
  3. 3.
    Liouville Evolution ($\rho_{k}$): The force on $\rho_{k}$ is nonlocal, mediated by the interpolation kernel $K(z)$ appearing in both the bulk quadrature points and the spectral coefficients of the boundary term:
    $\frac{d\rho_{k}}{d\tau}=-\frac{\partial S_{total}}{\partial\rho_{k}}+\nu_{\rho_{k}}$ (14.20)
  4. 4.
    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the exact algebraic constraints:
    $\frac{d\gamma_{1,k}}{d\tau}=-\mathrm{i}(\tilde{\xi}_{k}\xi_{k}-1)+\nu_{\gamma_{1,k}}$ (14.21)
In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the exact holomorphic gradient flow in complexified phase space. In Picard-Lefschetz theory, the union of all these deterministic flow lines originating from a complex saddle point constructs a middle-dimensional real manifold known as a Lefschetz thimble. If the noise $\nu$ were zero, the system would trivially collapse onto a 1D classical trajectory (the WKB approximation). However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, strictly incorporating all exact quantum fluctuations (the “vicinity” of the saddle points).
The phenomenon of exact S-matrix quantization we are looking for corresponds to a topological bifurcation of this stochastic system. At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift strictly vanishes along a continuous topological valley. Unimpeded by any restoring gradient force, the stochastic noise $\nu$ drives the complex trajectory to infinity along this flat zero mode. This infinite flat-valley integration geometrically generates the exact macroscopic S-matrix pole of the path integral, governed by Catastrophe Theory. Before studying this phenomenon in general form, we consider the simple solvable limit: the WKB quantization corresponding to the large rotational momentum of our rigid string.
@@ 15 The Linear Regge trajectories in the WKB limit
15.1 The Chiral Helicoid Saddle Point
In the semiclassical limit (large effective action), the path integral is dominated by the saddle point of the effective action. We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
@@ 15.1 The Chiral Helicoid Saddle Point
In the generic spinor formalism, with rescaling $w(z)=z^{-1/2}$, the conventional Helicoid is generated by $\mu(z)\propto(1,-z)^{T}$. However, as established in Appendix F, our geometric prescription requires enforcing the exact chiral (Hodge-dual) configuration. We freeze the right-handed twistor to a constant basis spinor:
$\lambda(z)=\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=\mathrm{i}\begin{pmatrix}1\\ 0\end{pmatrix}$ (15.2)
Substituting this into the null vector definition $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, we find that the spatial tangent vectors no longer identically vanish in the extra dimensions, like they did with conventional helicoid. Our Hodge dual minimal surface, with only one of two chiral Gauss maps (9.12), being active, and other being frozen, permanently breaks the 3D embedding constraint and allows the minimal surface to span the 4-dimensional space. Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid. The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
15.2 The Symmetrized Minkowski Helicoid
@@
The induced Minkowski metric for a generic string is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
Crucially, as derived in Appendix F, the conformal factor for the generic surface evaluates to $4\cos^{2}(\sigma_{M})$. In stark contrast, our chiral twistor evaluates to an induced Euclidean conformal factor $\exp\left(2\rho\right)\propto(1+e^{-2\sigma_{E}})$. Applying our physical prescription—taking the parity-even arithmetic mean of the Left-chiral (SD) and Right-chiral (ASD) configurations—upon double Wick rotation ($\sigma_{E}\to\mathrm{i}\sigma_{M}$), precisely isolates the real part:
$\text{Re}\left(1+e^{-2\mathrm{i}\sigma_{M}}\right)=1+\cos(2\sigma_{M})=2\cos^{2}(\sigma_{M})$ (15.4)
Thus, the effective proper area of the symmetrized chiral Minkowski Helicoid is exactly halved relative to the conventional generic string.
@@
Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space. Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity). By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space. Thus, the symmetrized Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
@@ 15.3 Semiclassical Motivation: Twisted Boundary Conditions
Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$. In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor444the perimeter renormalization terms in that factor are included in the physical quark mass.. This is relevant to our WKB approximation in the Wilson loop at large loops:
@@ 15.4 Effective Lagrangean
We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks. The confining factor introduces an effective Minkowski area term. We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m_{q}\int ds$. The total Lagrangian is:
@@
Because the arithmetic mean of the SD and ASD surfaces rigorously halves the classical area, the bulk contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is precisely halved relative to the generic string:
$L_{\text{helicoid}}$ $=-\frac{\sigma}{2}\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
$=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (15.9)
Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional. Adding the particle term $2m_{q}\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
$L(v,\omega)=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (15.10)
It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space. Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$. We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which, as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\sigma$.
@@
$J=\frac{\sigma}{2\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2m_{q}v^{2}}{\omega\sqrt{1-v^{2}}}$ (15.12)
@@
$E=\frac{\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (15.13)
@@ 15.5 Balance of quark forces and boundary condition
Recall that the total Lagrangian $L$ is composed of the bulk integral for the halved helicoid and the boundary terms for the quarks:
$L=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (15.14)
The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$. Differentiating the integral using the Leibniz rule yields the effective tension force, while differentiating the boundary term yields the centrifugal force:
$\frac{\partial L}{\partial R}=-\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (15.15)
Substituting $v=\omega R$ and computing the exact derivative of the mass term (properly applying the chain rule):
$-\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-v^{2}}}\right)=0$ (15.16)
@@
$\sigma\sqrt{1-v^{2}}=\frac{2m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+2m_{q}}$ (15.17)
This equation balances the physically halved confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
@@
These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$. To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$). In this limit, the mass terms vanish, and the system is governed purely by the symmetrized confining geometry:
$E\approx\frac{\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{2\omega^{2}}\frac{\pi}{2}$ (15.19)
@@
$J=\frac{\pi\sigma}{4}\frac{1}{\omega^{2}}=\frac{\pi\sigma}{4}\left(\frac{2E}{\pi\sigma}\right)^{2}\implies J=\frac{E^{2}}{\pi\sigma}$ (15.20)
@@
We thus flawlessly recover the exact linear Regge trajectory for the symmetrized chiral string:
$E^{2}=\pi\sigma J$ (15.21)